OpenStudy (anonymous):

The probability it will rain today is 0.4 and the probability it will rain tomorrow is 0.3. The probability it will rain both days is 0.2. What is the probability it will rain today or tomorrow?

OpenStudy (anonymous):

P(today or tomorrow) = P(today) + P(tomorrow) - P(today and tomorrow) = 0.4 + 0.3 - 0.2 = 0.5